# Sequential model confidence sets

### v1

@@ Appendix D SMCSs by FDR control
where the first inequality holds by convexity and the nonnegativity of $f_{t}$, the second inequality by the axioms of a norm, and the third inequality by the fact that $\boldsymbol{A},\boldsymbol{B}\in C_{t,1-\alpha}$. ∎
Appendix F Additional material for the Covid case study
F.1 Documentation of the models
The first forecasting model baseline is a naive model created by the Covid-19 forecast hub team as a comparison benchmark. Baseline simply outputs the outcome of the most recent week as a weekly median forecast and forms a predictive distribution around this median prediction based on the past weekly incidences. For more information, we refer to the supplementary material of Cramer et al. (2022b). Secondly, we consider the model MOBS-gleam, a global epidemic and mobility structured meta-population model (GLEaM), developed by the MOBS lab. It consists of a population-mobility layer, which uses transportation data to model the population mobility though a stochastic process, and of a disease model, which classifies each person in a population as susceptible, latent, infectious, or recovered, and estimates the rate in which a susceptible individual contracts the disease. For more details, we refer to the original MOBS lab publications by Balcan et al. (2009) and Balcan et al. (2010). The third model PSI-draft was developed by the Predictive Science Inc. It is a SEIRX model and estimates the reproductive number $R_{0}$ by classifying each individual of a population into susceptible (S), exposed (E), infected (I), removed (R), or quarantined (X). The model stratifies the population by age and disease severity, and incorporates the social restrictions imposed over the course of the pandemic, see the GitHub page of the Covid-19 forecast hub (https://github.com/reichlab/covid19-forecast-hub/tree/master/data-processed/PSI-DRAFT) or the Zoltar forecast archive (https://zoltardata.com/model/254) for more information. Further, we consider the neural network based model GT-deep, which includes some auto-regressive terms to ensure temporal consistency across consecutive predictions. GT-deep was the first purely data-driven model to be included in the Covid-19 forecast hub, in the sense that it makes no epidemiological assumptions about the spread of the pandemic. More detailed information can be found in the original publication by Rodriguez et al. (2021). Finally, we consider two models that made their predictions based on previously issued forecasts from other models: Firstly, ensemble, formed by the Covid-19 forecast hub team, which combines the latest predictions of all models that submitted a full set of forecasts at all $23$ quantile values. Whereas ensemble outputted the average forecast of all individual models at each quantile level for the first few weeks of the pandemic, it switched to the corresponding median forecast later on. As discussed in Cramer et al. (2022b), ensemble was the model used by the CDC in its public communications about the pandemic. Finally, CDC-ensemble, developed jointly by the Covid-19 forecast hub and CDC, is very similar to ensemble, but instead of the median of all models (at each quantile value), it considers only the 10 best models, measured by the average weighted interval score over the 12 most recent weeks, see Cramer et al. (2022a).
F.2 Comparison to the naive betting approach
Figure 7 shows the resulting SMCSs with respect to the naive approach, where $K_{ij,t}=2$ for all $i,j=1,\dots,m$ and $t\in\mathbb{N}$, compare to Section 5.1.2.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 7: The resulting SMCSs with respect to the naive choice $K_{ij,t}=2$ for all $i,j\leq m$ and $t\in\mathbb{N}$ with confidence level $\alpha=0.1$. Whereas, there are no significant differences to the adaptive betting scheme proposed in Section 5.1.2 for $\tau=0.3,0.5,0.9$, the adaptive approach is clearly superior for $\tau=0.1$.
Appendix G Station categories for the wind gust case study
Figure 8: Spatial distribution of the station categories over Germany.

### v3

@@ Appendix D SMCSs by FDR control
where the first inequality holds by convexity and the nonnegativity of $M_{t}$, the second one by the axioms of a norm, and the third one by the fact that $\boldsymbol{A},\boldsymbol{B}\in C_{t,1-\alpha}$. ∎
Appendix F Further simulation results
F.1 Uniformly weakly superior objects for unbounded losses
In Supplement B, it is shown that the score differences for quantile forecasts and probabilistic forecasts evaluated with the CRPS are conditionally bounded. In contrast, the highly relevant mean squared error (MSE) for mean forecasts does not satisfy this property. In the following, we present a simulation example where the MSE differences are conditionally sub-exponential. By using a small number of models, we enable graphical monitoring of the e-processes as measure of evidence over time.
For some data $(Y_{t})_{t\in\mathbb{N}}$, we consider $m=9$ mean forecasters issuing mean-forecasts
$m_{i,t}=Y_{t}+X_{i,t},\quad t\in\mathbb{N},$
with i.i.d. normally distributed errors $X_{i,t}\sim\mathcal{N}(\varepsilon_{i,t},1+\delta_{i,t})$ independent of the the sequence $(Y_{t})_{t\in\mathbb{N}}$. Analogously to Simulation 2, we let $(\varepsilon_{i,t},\delta_{i,t})\in\{-0.5,0,0.5\}^{2}$ be fixed over time, with the only exception that the superior model $i_{0}$ with zero bias and lowest variance is worse on Sundays, more precisely, $\delta_{i_{0},t}=-0.5$ and $\varepsilon_{i_{0},t}=0.6\cdot\mathbbm{1}_{\{t\in 7\mathbb{N}\}}$. If we assess the mean predictions by the MSE, we obtain the following conditional expected score differences, independent of the data $(Y_{t})_{t\in\mathbb{N}}$,
$L_{ij,t}=X_{i,t}^{2}-X_{j,t}^{2},\quad\mu_{ij,t}=\mathbb{E}(d_{ij,t})=\varepsilon_{i,t}^{2}+\delta_{i,t}-(\varepsilon_{j,t}^{2}+\delta_{j,t}),\quad t\in\mathbb{N}.$
Recall that the square of a Gaussian random variable is sub-exponential, and the difference of sub-exponential random variables is again sub-exponential. Therefore, for all $i,j\in\mathcal{M}_{0}$, the loss differences $(d_{ij,t})_{t\in\mathbb{N}}$ are sub-exponential with constants $\nu_{ij}=2(1+\delta_{i})+2(1+\delta_{j})$ and $\alpha_{ij}=4(1+\max\{\delta_{i},\delta_{j}\})$, that is, for all $t\in\mathbb{N}$,
$\mathbb{E}\left(e^{\lambda(d_{ij,t}-\mu_{ij,t})-\nu_{ij}^{2}\lambda^{2}/2}\big{|}\mathcal{F}_{t-1}\right)\leq 1,\quad|\lambda|<1/\alpha_{ij}.$
It follows that the cumulative product of the increments on the left-hand side is a non-negative supermartingale, and thus
$E_{ij,t}=\exp\left\{\lambda_{ij}t\hat{\Delta}_{ij,t}-\frac{t}{2}\nu_{ij}^{2}\lambda_{ij}^{2}\right\},\quad t\in\mathbb{N},$ (15)
is an e-process for $\mathcal{H}_{ij}^{\textrm{uw}}$, for any $0\leq\lambda_{ij}<1/\alpha_{ij}$. Note that the e-process at (15) could equivalently be expressed as in Proposition 3.3 with the variance process $V_{ij,t}=t\nu_{ij}$, see Howard et al. (2020, Proposition 5, Appendix E) for a discussion of the two equivalent notions of the sub-exponentiality property. We construct an SMCS for the uniformly weakly superior objects as given in Section 3.2 for $\alpha=0.1$. Figure 7 shows the average size of the SMCS over time with the averaged e-processes given in Figure 8.
Figure 7: Size of the SMCS over time, averaged over $N=1000$ simulations. The SMCS never wrongly rejects the superior model $i_{0}$ and contains $4.06$ models on average at the end of the evaluation period.
Model $\varepsilon$ $\delta$
1 $-0.5$ $-0.5$
2 $0.0/0.6$ $-0.5$
3 $0.5$ $-0.5$
4 $-0.5$ $0.0$
5 $0.0$ $0.0$
6 $0.5$ $0.0$
7 $-0.5$ $0.5$
8 $0.0$ $0.5$
9 $0.5$ $0.5$
Figure 8: Left: The averaged adjusted e-processes $(E_{i\cdot,t}^{\star})_{t\in\mathbb{N}}$ on a logarithmic scale for the models $i=1,\dots,9$ with the critical value $\log(1/\alpha)\approx 2.3$ given as a horizontal line. Right: Parameters of the models with superior model given by $i_{0}=2$.
F.2 Strongly superior objects for non-Gaussian data
Simulations 1 and 2 in the main text showcase the ability of sequential model confidence sets to deal with sequences of dependent data, while preserving anytime-validity. In these simulations, our sequential methods also exhibit remarkable statistical power. As shown in Figure 1, the original model set, which contains 49 different models, is narrowed down to less than 10 models after 1000 time steps. The data points are sampled from Gaussian distributions. One possible concern is that this design is not challenging enough, so the estimated statistical power might be overly optimistic.
In this section, we provide some empirical results on the performance of our methods in a non-Gaussian setting. We also compare them with the original model confidence set (MCS) of Hansen et al. (2011). Even though the original MCS is not anytime-valid, it is still be interesting to compare its size to our anytime-valid SMCSs. The simulation design is similar to the one of Simulation 1: we set $Y_{0}=0$ for convenience, and, for $n=1000$, we sample $(Y_{t})_{t=1}^{n}$ from the mixture of the Gaussian distributions
$\mathcal{N}\left(\arctan(Y_{t-1}),1+\sqrt{|Y_{t-1}|}\right),\quad\mathcal{N}\left(-\arctan(Y_{t-1}),1+\sqrt{|Y_{t-1}|}\right)$
with equal weights. The use of $\arctan(\cdot)$ prevents against violent oscillations over different orders of magnitude, which complicates the computation of the predictable bounds $c_{ij,t}$, and is rare in practice. The forecasting models are defined by introducing bias and dispersion parameters $\varepsilon,\delta$, in the same spirit with Simulations 1 and 2: at each step $t\geq 1$, forecaster $i$ outputs the mixture $F_{i,t}$ of the Gaussian distributions
$\mathcal{N}\left(\arctan(Y_{t-1})+\varepsilon_{i},1+\sqrt{|Y_{t-1}|}+\delta_{i}\right),\quad\mathcal{N}\left(-\arctan(Y_{t-1})+\varepsilon_{i},1+\sqrt{|Y_{t-1}|}+\delta_{i}\right),$
where $(\varepsilon,\delta)$ ranges over the set $\{-0.6,-0.4,\ldots,0.4,0.6\}^{2}$. The CRPS loss at a point $y$, for a Gaussian mixture $F$ with equally weighted components $\mathcal{N}(\mu_{1},\sigma_{1}^{2})$ and $\mathcal{N}(\mu_{2},\sigma_{2}^{2})$ is given by the following formula (Jordan et al., 2019):
$\frac{1}{2}A(y-\mu_{1},\sigma_{1}^{2})+\frac{1}{2}A(y-\mu_{2},\sigma_{2}^{2})-\frac{1}{2}A(0,2\sigma_{1}^{2})-\frac{1}{2}A(0,2\sigma_{2}^{2})-\frac{1}{2}A(\mu_{1}-\mu_{2},\sigma_{1}^{2}+\sigma_{2}^{2}),$
where $A(\mu,\sigma^{2})=\mu\left(2\Phi(\mu/\sigma)-1\right)+2\sigma\cdot\varphi(\mu/\sigma)$, and $\Phi,\varphi$ denote the CDF and the density of the standard normal distribution, respectively.
In contrast to Simulation 1, $\mu_{ij,t}=\mathbb{E}\left[d_{ij,t}\mid\mathcal{F}_{t-1}\right]$ now depends on $t$, so the relative performance of the different models can vary with time. This property makes this setting more realistic, but it raises an important question about the applicability of the MCS of Hansen et al. (2011). In particular, the theory in that paper is developed around the assumption that the unconditional mean $\mathbb{E}[d_{ij,t}]$ does not depend on $t$. A simulation of $10000$ trajectories of the sequence $(Y_{t})_{t=1}^{1000}$ and the corresponding forecasts $F_{i,t}$ confirms that $\mathbb{E}[L_{i,t}]$ does not have any clear temporal trends, see Figure 9. Therefore, the assumption of a time-independent expected loss is approximately satisfied.
Figure 9: Expected CRPS of the forecaster that corresponds to $(\varepsilon,\delta)=(-0.4,-0.6)$.
The MCS procedure is applied sequentially in the following way: We start with the original set $\mathcal{M}_{0}=\widehat{\mathcal{M}}_{0}$ containing all $49$ models. During the first $50$ time points we do not perform any statistical tests. Instead, we only observe the outcomes and the report the loss of each forecast. This warm-up period is necessary due to the Bootstrap estimation of the variance that the MCS algorithm makes use of. Afterwards, for each time point $t=51,52,\ldots,1000$:
  1. 1.
    We observe the $t$-th outcome $Y_{t}$.
  2. 2.
    We test the hypothesis that all remaining models have equal expected loss, namely,
    $H_{0,\widehat{\mathcal{M}}_{t-1}}:\mu_{ij}=0\text{ for all }i,j\in\widehat{\mathcal{M}}_{t-1},$
    where $\mu_{ij}=\mathbb{E}[d_{ij,t}]$ denotes the expected loss difference between models $i$ and $j$. As we mentioned earlier, it is assumed that $\mu_{ij}$ does not depend on $t$, so we have omitted the index $t$. We use the test statistic
    $T_{\text{R},\text{M}}=\max_{i,j\in\widehat{\mathcal{M}}_{t-1}}\left|t_{ij}\right|,\text{ where }t_{ij}=\frac{\overline{d}_{ij}}{\sqrt{\widehat{\text{var}}\left(\overline{d}_{ij}\right)}},$
    and $\overline{d}_{ij}=n^{-1}\sum_{s=1}^{t}d_{ij,s}$ is the average relative loss of models $i,j$. The quantity $\widehat{\text{var}}\left(\overline{d}_{ij}\right)$ is a Bootstrap estimate of $\text{var}\left(\overline{d}_{ij}\right)$ based on the first $t$ observed points.
  3. 3.
    If $H_{0,\widehat{\mathcal{M}}_{t-1}}$ is not rejected, then we set ^$\widehat{\mathcal{M}}_{t}=\widehat{\mathcal{M}}_{t-1}$ and $t\leftarrow t+1$, and we go back to step 1. Else, we eliminate the worst-performing model, namely, the one indexed by
    $e_{\text{R},\text{M}}=\operatorname*{arg\,max}_{i\in\widehat{\mathcal{M}}_{t-1}}\left\{\sup_{j\in\widehat{\mathcal{M}}_{t-1}}\frac{\overline{d}_{ij}}{\sqrt{\widehat{\text{var}}\left(\overline{d}_{ij}\right)}}\right\}.$
    We then set ^$\widehat{\mathcal{M}}_{t-1}\leftarrow\widehat{\mathcal{M}}_{t-1}\backslash\{e_{\text{R,M}}\}$ and go back to step 2.
For the variance estimation in step 2, we used $B=2000$ Bootstrap samples. For comparison, Hansen et al. (2011) used $B=1000$ samples in their simulations. In the documentation of the R-package MCS, Bernardi and Catania (2014) set $B=5000$ as the default. We tried this option, too, but the computational cost was too high, so we opted for a lower, but still reliable number of Boostrap iterations.
In the above implementation of the MCS algorithm, after a model is eliminated, it is totally discarded in future time points. That is, when we set $t\leftarrow t+1$ and move back to step 1, we only repeat the process for the models in ^$\widehat{\mathcal{M}}_{t}$, and not for all models in $\mathcal{M}_{0}$. For anytime-valid methods, this is not possible, as illustrated by Equation (6), where $m$ stays constant over time.
We implemented the sequential version of the MCS procedure like this, since it leads to lower computational cost. As explained earlier, the MCS algorithm involves the Bootstrap estimation $\widehat{\text{var}}\left(\overline{d}_{ij}\right)$ after every single model rejection. This is already expensive computationally, and the cost would be higher if we restarted with the full set of models at every new time point.
However, this variation has no impact on the rate of coverage of the optimal model: since we are testing the strong hypothesis, one rejection of this model is enough to exclude it from future confidence sets. Even though the losses associated with this model would otherwise be used in future time steps, the model itself would not reenter the model confidence set.
However, the size of the model confidence set could be affected. After the removal of a number of models, the definition of $\overline{d}_{ij}$ changes, and the Bootstrap estimation of the variance is based only on the surviving models. Even in this case, the size of the MCS would be pushed down, because it would be easier for sub-optimal models $i\in\mathcal{M}_{0}$ to get excluded in future steps. This happens because the excluded models $j$ that are likely worse than $i$ would otherwise push the quantiles of the distribution of $\overline{d}_{ij}$ upwards. Now that they are not taken into account, performance differences between $i$ and better-performing models are going to become more pronounced.
As in Figure 1, we track the average size of the confidence set and the rate of coverage of the optimal forecast across a period of $T=1000$ time points. The averages are computed over $N=1000$ Monte Carlo iterations and the results are shown in Figure 10.
As expected, the original MCS has very poor sequential control of the Type I error. In contrast, the anytime-valid SMCSs contains the superior model across the entire time domain. Our SMCS are uniformly larger than the (boosted) MCS run sequentially over time. However, this comes at the cost of very poor coverage rates.
Figure 10: Comparison of the original MCS with SMCS. Left: Coverage rates; Right: Average size of MCS and SMCS.